OpenStudy (ttop0816):

help please! On the interval 0 ≤ x ≤ 2 π , where are the asymptotes for the graph of y = cot(x - pi/2 ) located? 0, π and 2π π and 2π π/2 and 3π/2 << i chose this! π/2, π and 3π/2

OpenStudy (mathmale):

We could simplify this problem a little by asking, "Where is tan(x-pi/2) = 0 on [0,2pi]? That's the easiest way to identify the asymptotes of the given function, because when tan(x-pi/2)=0, cot(x-pi/2) is undefinted.

OpenStudy (anonymous):

Answer is c. You are right

OpenStudy (ttop0816):

@mathmale then would my answer be correct or incorrect...??

OpenStudy (anonymous):

Correct

OpenStudy (ttop0816):

@MayankD how about for this question if you dont mind! https://media.glynlyon.com/g_trg_2013/2/graph_sec(x-pi_2).gif y = csc(x - pi/2) <<< chose this! y = sec(x - pi/2)

OpenStudy (anonymous):

Its sec(x-pi/2)

OpenStudy (anonymous):

When in doubt substitute values and compare

OpenStudy (ttop0816):

@MayankD thank you so much! im correct for this too right..? im sorry if im bothering you! The graph of y = tanx + 1 crosses the y-axis at _____. -1 0 1<< 2

OpenStudy (anonymous):

Yes